OpenStudy (anonymous):

how do i evaluate cot pie over 7?

OpenStudy (amoodarya):

\[\frac{ \pi }{7 }\] is not a famous angle so you have to use calculator \[\cot \frac{ \pi }{7 } =\frac{\cos \frac{ \pi }{7 } }{\sin\frac{ \pi }{7 } }\]

OpenStudy (anonymous):

what does cot mean though?

OpenStudy (amoodarya):



OpenStudy (amoodarya):

\[\cot x= \\\frac{ 1 }{ \tan x }=\\\frac{ 1 }{ \frac{ sinx }{ \cos x } } =\\\frac{ \cos x }{ \sin x }\]